OpenStudy (anonymous):

A copy machine makes 105 copies in 3 minutes and 45 seconds. How many copies does it make per minute?

OpenStudy (shinalcantara):

remember that a minute is 60 sseconds then you'll have it as: \[\frac{ 105 }{ 3 + \frac{ 45 }{ 60 } } = \frac{ 105 }{ \frac{ 225 }{ 60 } }\] \[ \frac{ 105 }{ \frac{ 225 }{ 60 } } = \frac{ 60(105) }{ 225 }\] = 28pcs/min

OpenStudy (confluxepic):

Correct.